OpenStudy (anonymous):

Express :: 4 cos⁡θ+ 6 sin⁡θ in the form of R sin⁡(θ+α),where R>0 and 0<α<1/2 π please help :)

OpenStudy (cwrw238):

use the relation R sin⁡(θ+α) = Rsinθcosα + Rcosθsinα

OpenStudy (cwrw238):

compare coefficients of cosθ and sinθ with the RHS of the above relation for sin θ you will get Rcosα = 6 do you see how i got that?

OpenStudy (cwrw238):

and what would Rsinα equal to?

OpenStudy (anonymous):

what's an RHS ? unfortunately no :(

OpenStudy (cwrw238):

RHS = right hand side Rsinθcosα + Rcosθsinα Rcosα sinθ + Rsinα cosθ compare: 6 sinθ + 4 cosθ

OpenStudy (cwrw238):

so Rcosα = 6 and Rsinα = 4

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

yeah , understood :D

OpenStudy (cwrw238):

ok now from the above 2 equations we can find the value of R and α as follows Rsinα ------ = tanα = 2/3 - solve this to get α Rcosα find R R^2(sin^2α + cos^2α) = 6^2 + 4^2 R^2 = 100

OpenStudy (anonymous):

Why did you combine them by division ? The mark scheme values are different ..

OpenStudy (anonymous):

for α it's the same . but the R^2 it's = 4^2 + 6^2 which is weird

OpenStudy (cwrw238):

silly me 6^2 + 4^2 = 52 not 100!!!

OpenStudy (cwrw238):

i skipped a step R^2cos^2α + R^2sin^2α = 4^2 + 6^2 factoring R^2(cos^2α + sin^2α) = 4^2 + 6^2

OpenStudy (cwrw238):

combining by division eliminates R

OpenStudy (anonymous):

yeah noticed it now :D The second step , powering everything ..?

OpenStudy (cwrw238):

yes squaring each term and sin^2α + cos^2α = 1

OpenStudy (anonymous):

That's logically weird , Thanks so much and sorry such questions :D Appreciate it :)

OpenStudy (cwrw238):

yw